# RocqStar: Leveraging Similarity-driven Retrieval and Agentic Systems for Rocq generation

### v1

@@ Abstract
Interactive Theorem Proving was repeatedly shown to be fruitful combined with Generative Artificial Intelligence. This paper assesses multiple approaches to Rocq generation and illuminates potential avenues for improvement. We highlight the importance of thorough premise selection for generating Rocq proofs and propose a novel approach, leveraging retrieval via a self-attentive embedder model. The evaluation of the designed approach shows up to 28% relative increase of the generator’s performance. We tackle the problem of writing Rocq proofs using a multi-stage agentic system, tailored for formal verification, and demonstrate its high effectiveness. We conduct an ablation study and show the use of multi-agent debate on the planning stage of proof synthesis.
@@ 1 Introduction
In recent years, the advent of Generative Artificial Intelligence (AI) has accelerated the process of developing new software. However, there are studies [21] showing that users who use AI assistants tend to introduce more bugs and vulnerabilities into their code, compared to those who write code on their own. Formal software verification could help mitigate the issue of bugs and security flaws, as it ensures that the software operates correctly and reliably in compliance with the given specification. Under the assumption of a well-formed specification, formal verification provides strong guarantees and an acceptance criterion for the generated code. Interactive Theorem Prover (ITP) is a software tool that assists the user with the development of formal specifications and proofs. To date, there exist several ITPs, such as Rocq (former Coq) [2], Lean [5], Agda [13], Isabelle [19], and others. Rocq is a mature ITP, which has experienced more than 30 years of continuous development and improvement. Rocq has an extensive track record of high-impact projects. For example, Rocq was used to verify the correctness of the CompCert C compiler [15], the only compiler, in which an extensive study found no bugs [33].
@@
Verifying software has always been a rigorous and time-consuming process requiring much human effort. A number of solutions have been developed to help automate the process of theorem proving in Rocq. Proofs in Rocq are constructed from so-called tactics, which are elementary building blocks. Using tactics, the user manipulates the proof state — a data structure, which contains the current goal and the context of the proof. Thus, with every applied tactic, the task is transformed and could be solved recursively. Most solutions implement tactic-prediction approaches and employ beam search or a similar algorithm to navigate the search space. Tactician [4] is a KNN-based approach, which does similarity-based retrieval of tactics used in similar states. CoqGym [31] and Proverbot9001 [25] use Recurrent neural networks, Graph2Tac [24] proposed a novel graph-based neural tactic prediction. Thakur et al. [26] and Kozyrev et al. [14] instead build generation pipelines around general-purpose, cloud-hosted LLMs, so that no heavy computations occur on the user’s machine. CoqPilot [14], along with that, contributes a benchmarking framework and allows seamless integration of standalone tools into the workflow of Rocq’s user.
@@
Many approaches call attention to the importance of premise selection, i.e., retrieving useful context information to advance generation. Yang et al. [32] introduced LeanDojo, a retrieval-augmented prover in Lean that significantly improves over non-retrieval baselines. Thompson et al. [27] present the Rango tool and report state-of-the-art performance on the CoqStoq benchmark, automatically synthesizing complete proofs for 32% of the theorems. The work highlights how strongly the well-formed context contributes to the success of Rango. Moreover, they show that proof retrieval is the most performant mechanism for premise selection. The proof retriever selects relevant previously completed proofs from the current project and provides them as references to the model. According to the evaluation, Rango proved 47% more theorems than the variant without a proof retriever. However, their mechanism of retrieving proofs relies on the baseline text similarity over states. In this work, we build on top of their research and propose a novel embedding model for Rocq statements. It is trained to predict the similarity of their proofs and shows relative improvement of up to 28% on the evaluation set.
@@
Another promising direction in generative theorem proving that we have identified is Agentic Systems. Research by Kozyrev et al. [14] shows that current Rocq generation methods mostly struggle with complex reasoning tasks. Algorithms that perform proof search on top of a tactic generator slow down dramatically and suffer performance degradation as theorem complexity grows, due to the properties of tree-based search. Other neural methods, which apply LLMs, suffer from the same problem due to the inability of the model to handle complex reasoning tasks [11]. Agentic systems are known to address these problems; however, to our knowledge, there were close to no attempts to build an autonomous agentic system for an ITP. We build an extensive Model Context Protocol (MCP) server for Rocq and implement an autonomous Agentic System over it, utilizing various problem-specific solutions, such as multi-agent debate. We conduct an evaluation and show that our agentic system strongly outperforms all other previously benchmarked solutions in the CoqPilot’ work, raising the ratio of successfully proven theorems from 51% to 60%.
@@ 1.1 Contributions
The main contributions of this paper are:
@@
RocqStar proof retriever  We propose a novel approach for premise selection in Rocq. Rocq suffers from the data-scarcity problem that is common to most ITPs. Aggregating the largest publicly available repositories, one could expect to collect roughly 300 million tokens of Rocq, and about the same for Lean. In contrast, open-source Python corpora easily exceed 100 billion tokens. To tackle this issue we contribute a convenient standalone tool BigRocq to extract additional data from Rocq code, utilizing the nature of Rocq’s system and the intermediate states of the proof. BigRocq bridges the gap between Automated Generation and Rocq’s ecosystem. Using BigRocq, we mine a dataset of 76,524 statements with corresponding proofs from 4 big projects and train a self-attentive embedder model, which learns to predict how close the proofs of given statements will be. In addition, we provide a pipeline to reproduce such embeddings for an arbitrary project, which offers even better results. We integrate the solution as a new retrieval approach for selecting context theorems in CoqPilot and evaluate it using CoqPilot’s benchmarking infrastructure. Compared to the baseline text similarity-based ranker, we show an improvement of 28% on the evaluation set. BigRocq tool, training dataset, and the code for training the embedder model are available at https://github.com/JetBrains-Research/rocqstar-rag. A model checkpoint is available at https://huggingface.co/JetBrains-Research/rocq-language-theorem-embeddings.
RocqStar agentic system  Addressing the lack of research of applying agentic systems to ITPs, we build an autonomous Agent for writing Rocq proofs. A custom MCP server built over coq-lsp [6] handles interaction with Rocq, its code is available at https://github.com/JetBrains-Research/rocqstar-agentic-system. We implement an agentic system that includes such stages as planning, execution, and reflection. An ablation study demonstrates the critical role of planning, particularly the multi-agent debates (MAD) framework, in boosting performance. Evaluation shows that our end-to-end agent can solve 60% of theorems from the CoqPilot’s dataset. To deploy our AI Agent, we use privately available infrastructure called IDEFormer, but all our agent’s code is available at https://github.com/JetBrains-Research/rocqstar-agentic-system/tree/main/rocqstar-agent.
@@
The remainder of the paper is organized as follows. § 2 describes our Similarity-Driven Retrieval mechanism and § 3 introduces the agentic system. The retrieval component is evaluated in § 4.1 and the agent in § 4.2. § 4.3 provides an ablation study of the agentic system. We describe the related work in § 5 and conclude in § 6.
@@ 2 Similarity-driven Retrieval
We distinguish two ways of doing premise selection in Rocq. Hint selection — given a context $C$ and a tactic with an unknown positional argument, e.g. apply _, the task is to yield potential candidates for the argument. Proof selection, in turn, given theorem statement $S$, focuses on choosing other statements with their respective proofs, so that their presence in the context of the generation request would help the model with the generation of the proof for statement $S$. Most works [3, 12, 27, 32] on premise selection in Rocq and other ITPs focused on doing hint selection. However, Thompson et al. [27] and Kozyrev et al. [14] show that even a baseline proof selection significantly boosts the model’s capabilities and is stronger than hint selection. The baseline proof selection presented in both works [27, 14], given the target statement $s_{*}$ and a database of already proven theorems $\left[s_{0},p_{0}\right],\dots,\left[s_{n},p_{n}\right]$, chooses theorems, statements of which have the maximum similarity to the target one. Similarity is defined by the BM-25 information retrieval technique [23] or Jaccard similarity index.
Both existing approaches suppose that if statements $s_{*}$ and $s_{i}$ are similar, their respective proofs $p_{*}$ and $p_{i}$ are similar as well:
$\operatorname{similarity(s_{*},s_{i})}\Longrightarrow\operatorname{similarity(p_{*},p_{i})}$
therefore assume that theorems $\{\left[s_{j},p_{j}\right]\}$, chosen in such a manner, are relevant while proving $s_{*}$. However, we show that this implication often does not hold. Let us define the proof similarity $D_{L}$ as the Levenshtein edit distance on lists of tactics, where the cost of substitution between two tactics is the Levenshtein distance over their strings. We include a Jaccard similarity term and add noise for robustness; otherwise, the proof-distance distribution over randomly selected pairs of theorems becomes U-shaped and the model fails to learn.
$p_{i}=\left[tac_{i_{0}},\dots,tac_{i_{m}}\right],\quad l_{i}=|s_{i}|,\quad D_{L}(p_{i},p_{j})=\frac{\mathrm{Lev}(p_{i},p_{j})}{\max(l_{i},l_{j})},\quad D_{J}(p_{i},p_{j})=1-\frac{|p_{i}\cap p_{j}|}{|p_{i}\cup p_{j}|}$
$\mathrm{proof\_distance}(p_{i},p_{j})=\alpha\,D_{L}(p_{i},p_{j})+(1-\alpha)\,D_{J}(p_{i},p_{j})+\gamma,\quad\alpha=0.7,\quad\gamma\sim\mathcal{U}(-\varepsilon,+\varepsilon)$
@@
Considering 1,855,701 pairs of theorems from the IMM project [^1] IMM  https://github.com/weakmemory/imm, we compute correlations between statement similarities and respective proof similarities. In summary, BM25-based statement similarity shows a weak negative relationship with Levenshtein-based proof distance (Pearson $r$ = –0.154, Spearman $\rho$ = –0.171). In contrast, BM25-based proof similarity exhibits near-zero Pearson correlation ($r=0.029$) and a small positive Spearman correlation ($\rho=0.240$) — in both cases, effectively negligible.
@@
To assess the issue of ineffective proof selection, we try to find such function $f(s_{i},s_{j})$ that correlates with $\operatorname{proof\_distance}(\operatorname{proof}_{i},\operatorname{proof}_{j})$ stronger than statement similarity. In this work, we introduce a neural method that learns vector embeddings for Rocq theorem statements, training them so that the distance between any two vectors mirrors the similarity between the theorems’ proofs.
@@ 2.1 Dataset mining
Along with other ITPs, Rocq struggles with data scarcity. To assess this issue, we mine additional data from the Rocq code. We utilize Rocq system’s functionality, preprocess theorems, and transform sequential proof structures into trees. Fig. 1 illustrates an example of the process. Since every node in such tree is a valid state, we can automatically construct a proof for it, recursively iterating through its subtree edges. By extracting the statements with corresponding proofs, we can enlarge an arbitrary dataset of Rocq theorems roughly by a magnitude of 4. Dataset format and its details are described in Appendix B.
Theorem test : forall n : nat, n = 0 \/ n <> 0. Proof. intros n. destruct n. - left; auto. - right; auto. Qed. ${\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{2}}$ []${\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{0}}$${\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{1}}$${\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{3}}$ []left; auto.intros n.destruct n.destruct n.right; auto. Lemma s1 (n : nat) : n = 0 \/ n <> 0. Proof. destruct n. - left; auto. - right; auto. Qed. Lemma s2 : 0 = 0 \/ 0 <> 0. Proof. left; auto. Qed. Lemma s3 (n : nat) : S n = 0 \/ S n <> 0. Proof. right; auto. Qed.
@@
We call the proposed tool BigRocq and make it publicly available as a standalone component of our system. The idea of mining additional training data from the intermediate states of the ITP is not new; Kogkalidis et al. [12] conducted analogous research for the Agda [13] language. Similar research for Rocq also takes place; however, some of those works are highly dependent on the deprecated ways of communication with Rocq’s compiler [31] and do not support up-to-date versions of Rocq. In contrast, others implement similar ideas as a part of the training pipeline and do not allow for seamless reuse. Using BigRocq, we mine a total of 76,524 statements, collected from 344 files from 4 big Rocq projects.

### v2

# The 4th Deep Learning for Code Workshop

@@ Abstract
Interactive Theorem Proving was repeatedly shown to be fruitful combined with Generative Artificial Intelligence. This paper assesses multiple approaches to Rocq generation and illuminates potential avenues for improvement. We highlight the importance of thorough premise selection for generating Rocq proofs and propose a novel approach, leveraging retrieval via a self-attentive embedder model. The evaluation of the designed approach shows up to 28% relative increase of the generator’s performance. We tackle the problem of writing Rocq proofs using a multi-stage agentic system, tailored for formal verification, and demonstrate its high effectiveness. We conduct an ablation study and demonstrate shows that incorporating multi-agent debate during the planning stage increases the proof success rate by 20% overall and nearly doubles it for complex theorems, while the reflection mechanism further enhances stability and consistency.
@@ 1 Introduction
In recent years, the advent of Generative Artificial Intelligence (AI) has accelerated the process of developing new software. However, there are studies [22] showing that users who use AI assistants tend to introduce more bugs and vulnerabilities into their code, compared to those who write code on their own. Formal software verification could help mitigate the issue of bugs and security flaws, as it ensures that the software operates correctly and reliably in compliance with the given specification. Under the assumption of a well-formed specification, formal verification provides strong guarantees and an acceptance criterion for the generated code. Interactive Theorem Prover (ITP) is a software tool that assists the user with the development of formal specifications and proofs. To date, there exist several ITPs, such as Rocq (former Coq) [2], Lean [6], Agda [14], Isabelle [20], and others. Rocq is a mature ITP, which has experienced more than 30 years of continuous development and improvement. Rocq has an extensive track record of high-impact projects. For example, Rocq was used to verify the correctness of the CompCert C compiler [16], the only C compiler, in which an extensive study found no bugs [33].
@@
Verifying software has always been a rigorous process requiring much time and human effort. A number of solutions have been developed to help automate the process of theorem proving in Rocq. Proofs in Rocq are constructed from so-called tactics, which are elementary building blocks. Using tactics, the user manipulates the proof state — a data structure, which contains the current goal and the context of the proof. Thus, with every applied tactic, the task is transformed and could be solved recursively. Most solutions implement tactic-prediction approaches and employ beam search or a similar algorithm to navigate the search space. Tactician [4] is a KNN-based approach, which does similarity-based retrieval of tactics used in similar states. CoqGym [31] and Proverbot9001 [26] use Recurrent Neural Networks (RNNs), Graph2Tac [25] proposed a novel graph-based neural tactic prediction. Thakur et al. [27] and Kozyrev et al. [15] instead build generation pipelines around general-purpose, cloud-hosted LLMs, so that no heavy computations occur on the user’s machine. CoqPilot [15], along with that, contributes a benchmarking framework and allows seamless integration of standalone tools into the workflow of Rocq’s user.
@@
Many approaches call attention to the importance of premise selection, i.e., retrieving useful context information to advance generation. Yang et al. [32] introduced LeanDojo, a retrieval-augmented prover in Lean that significantly improves over non-retrieval baselines. Thompson et al. [28] present the Rango tool and report state-of-the-art performance on the CoqStoq benchmark, automatically synthesizing complete proofs for 32% of the theorems. The work highlights how strongly the well-formed context contributes to the success of Rango. Moreover, they show that proof retrieval is the most performant mechanism for premise selection. The proof retriever selects relevant previously completed proofs from the current project and provides them as references to the model. According to the evaluation, Rango proved 47% more theorems than the variant without a proof retriever. However, their retrieval mechanism assumes that two textually similar statements have proofs relevant to each other. In this work, we demonstrate that this assumption oversimplifies the relationship between statements and proofs and introduce a novel embedding model for Rocq statements. It is trained to predict the similarity between their proofs and achieves up to a 28% relative improvement on the evaluation set.
@@
Another promising direction in generative theorem proving that we have identified is Agentic Systems. Research by Kozyrev et al. [15] shows that current Rocq generation methods mostly struggle with complex reasoning tasks. Algorithms that perform proof search on top of a tactic generator slow down dramatically and suffer performance degradation as theorem complexity grows, due to the properties of tree-based search. Other neural methods, which apply LLMs, suffer from the same problem due to the inability of the model to handle complex reasoning tasks [12]. Agentic systems are known to address these problems; however, to our knowledge, there were close to no attempts to build an autonomous agentic system for an ITP. We build an extensive Model Context Protocol (MCP) server for Rocq and implement an autonomous Agentic System over it, utilizing various problem-specific solutions, such as multi-agent debate. We conduct an evaluation and show that our agentic system strongly outperforms all other previously benchmarked solutions in the CoqPilot’s work, raising the ratio of successfully proven theorems from 51% to 60%.
@@ 1.1 Contributions
The main contributions of this paper are the following.
@@
RocqStar proof retriever  We propose a novel approach for premise selection in Rocq. Rocq suffers from the data-scarcity problem that is common to most ITPs. Aggregating the largest publicly available repositories, one could expect to collect roughly 300 million tokens of Rocq, and about the same for Lean. In contrast, open-source Python corpora easily exceed 100 billion tokens. To tackle this issue we contribute a convenient standalone tool BigRocq to extract additional data from Rocq code, utilizing the nature of Rocq’s system and the intermediate states of the proof. BigRocq bridges the gap between Automated Generation and Rocq’s ecosystem. Using BigRocq, we mine a dataset of 76,524 statements with corresponding proofs from 4 big projects and train a self-attentive embedder model, which learns to predict how close the proofs of given statements will be. In addition, we provide a pipeline to reproduce such embeddings for an arbitrary project, which offers even better results. We integrate the solution as a new retrieval approach for selecting context theorems in CoqPilot and evaluate it using CoqPilot’s benchmarking infrastructure. Compared to the baseline text similarity-based ranker, we show an improvement of 28% on the evaluation set. The BigRocq tool, the training dataset, and the code for training the embedder model are available at https://github.com/JetBrains-Research/rocqstar-rag. The embedder model’s checkpoint is available at https://huggingface.co/JetBrains-Research/rocq-language-theorem-embeddings.
RocqStar agentic system  Addressing the lack of research on applying agentic systems to ITPs, we build an autonomous system for generating Rocq proofs. A custom MCP server built on top of coq-lsp [7] handles the interaction with Rocq; its source code is available at https://github.com/JetBrains-Research/rocqstar-agentic-system/tree/main/mcpServer. Our approach follows a structured process consisting of planning, execution, and reflection. An ablation study shows that while naive planning has limited impact, effective planning based on the Multi-Agent Debate (MAD) framework plays a crucial role. Specifically, it yields a 20% relative improvement in the overall proof success rate and nearly doubles the success rate on complex theorems with longer reference proofs (33% vs. 17%). Additionally, we demonstrate the benefits of the reflection mechanism, which improves the overall proof success rate from 48% to 66% and more than quadruples success on complex theorems; see §˜4.3 for details. The evaluation results show that the RocqStar system solves up to 60% of theorems from the CoqPilot dataset. It is implemented using Koog [^1] Koog  https://docs.koog.ai, an innovative framework for building AI agents by JetBrains. The source code is available at https://github.com/JetBrains-Research/rocqstar-agentic-system/tree/main/koogAgent.
@@
The remainder of the paper is organized as follows. §˜2 describes our Similarity-Driven Retrieval mechanism. §˜3 introduces the agentic system. §˜4 presents an evaluation of the retrieval component (§˜4.1), the agent (§˜4.2), and an ablation study of the agentic system (§˜4.3). We describe the related work in §˜5 and conclude in §˜6.
@@ 2 Similarity-driven Retrieval
We distinguish two ways of doing premise selection in Rocq. Hint selection — given a context $C$ and a tactic with an unknown positional argument, e.g. apply _, the task is to yield potential candidates for the argument. Proof selection, in turn, given theorem statement $S$, focuses on choosing other statements with their respective proofs, so that their presence in the context of the generation request would help the model with the generation of the proof for statement $S$. Since the approach of applying general purpose models to proof generation is relatively new, most of the works [3, 13, 28, 32] on premise selection in Rocq and other ITPs focused on hint selection. However, Thompson et al. [28] and Kozyrev et al. [15] show that even a baseline proof selection significantly boosts the model’s capabilities and is stronger than hint selection. The baseline proof selection presented in both works [28, 15], given the target statement $s_{*}$ and a database of already proven theorems $\left[s_{0},p_{0}\right],\dots,\left[s_{n},p_{n}\right]$, chooses theorems, statements of which have the maximum similarity to the target one. Similarity is defined by the BM-25 information retrieval technique [24] or the Jaccard similarity index. That results in packing the generator’s context with theorems, based on how similar their statements are syntactically.
We propose a retrieval mechanism that improves the performance of the generator compared to the described baseline. During the generation of the target proof for the statement $S$ we generally assume that the model benefits more from seeing similar proofs to the one it needs to generate, rather than from seeing similar statements with proofs dissimilar from the target one. Evaluation of our retriever in §˜4.1 supports this supposition. One might assume that if statements $s_{*}$ and $s_{i}$ are similar, their respective proofs $p_{*}$ and $p_{i}$ are similar as well:
$\operatorname{similarity}(s_{*},s_{i})\;\Longrightarrow\;\operatorname{similarity}(p_{*},p_{i})$ (1)
However, we show that this implication often does not hold. The heuristic of retrieving similar statements produces decent baseline results, but fails in complex cases, leaving room for improvement. We design our retrieval method to guide context selection based on the similarity of the proofs and show its practicality.
Let us define the proof similarity $D_{L}$ as the Levenshtein edit distance computed over lists of tactics. Insertions and deletions of tactics have a unit cost, as in the standard Levenshtein formulation. The substitution cost between two tactics is proportional to the Levenshtein distance between their string representations. The resulting distance is normalized by the maximum proof length.
$p_{i}=\left[tac_{i_{0}},\dots,tac_{i_{m}}\right],\quad l_{i}=|s_{i}|,\quad D_{L}(p_{i},p_{j})=\frac{\mathrm{Lev}(p_{i},p_{j})}{\max(l_{i},l_{j})}$
@@
We conduct the following experiment to examine whether the relation in Equation˜1 holds in practice. Considering 1,855,701 pairs of theorems from the IMM project [^2] IMM  https://github.com/weakmemory/imm, we compute correlations between statement similarities and respective proof similarities. In summary, BM25-based statement similarity shows a weak negative relationship with the Levenshtein-based proof distance (Pearson $r$ = –0.154, Spearman $\rho$ = –0.171). The code to reproduce these experiments could be found in the RocqStar-retriever repository [^3] RocqStar retriever: https://github.com/JetBrains-Research/rocqstar-rag/tree/main/experiments.
@@
To assess the issue of ineffective proof selection, we try to find a function $f(s_{i},s_{j})$ that correlates with the defined proof-distance stronger than statement similarity does. In this work, we introduce a neural method that learns vector embeddings for Rocq theorem statements, training them so that the distance between any two vectors mirrors the similarity between the respective theorem’ proofs.
@@ 2.1 Dataset mining
Along with other ITPs, Rocq struggles with data scarcity. To assess this issue, we mine additional data from the Rocq code. We utilize Rocq system’s functionality, preprocess theorems, and transform sequential proof structures into trees. Fig.˜1 illustrates this transformation process. Since every node in such a tree is a valid state, we can automatically construct a proof for it by recursively following its subtree edges. Extracting all intermediate statements together with their proofs allows us to expand any given Rocq theorem dataset by a factor roughly proportional to the average proof length. In our case, this resulted in an approximately fourfold increase in dataset size. The dataset format and further details are provided in Appendix B.
Theorem test : forall n : nat, n = 0 \/ n <> 0. Proof. intros n. destruct n. - left; auto. - right; auto. Qed. $\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{2}$} $\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{0}$$\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{1}$$\color[rgb]{0.81640625,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.81640625,0,0}s_{3}$} []left; auto.intros n.destruct n.destruct n.right; auto. Lemma s1 (n : nat) : n = 0 \/ n <> 0. Proof. destruct n. - left; auto. - right; auto. Qed. Lemma s2 : 0 = 0 \/ 0 <> 0. Proof. left; auto. Qed. Lemma s3 (n : nat) : S n = 0 \/ S n <> 0. Proof. right; auto. Qed.
@@
We call the proposed tool BigRocq and make it publicly available as a standalone component of our system. The idea of mining additional training data from the intermediate states of the ITP is not new; Kogkalidis et al. [13] conducted analogous research for the Agda [14] language. Similar research for Rocq also takes place [31, 25]; however, some of those works are highly dependent on the deprecated ways of communication with Rocq’s compiler [31] and do not support up-to-date versions of Rocq. In contrast, others implement similar ideas as a part of the training pipeline and do not allow for seamless reuse. Using BigRocq, we mine a total of 76,524 statements, collected from 344 files from 4 big Rocq projects.